OpenStudy (anonymous):

in science the percent error formula is calculated using |a-p|/p x 100=e where a is the actual amount, p is the predicted amount, and e is the percent error. Find the actual amount if the predicted amount is 10 grams and percent error is 21.5 a. 7.85 b. 178.5 c.10.215 d.9.785

OpenStudy (ash2326):

@mathdummy4 we have \[\frac{|a-p|}{p}\times 100=e\] we have p=10 grams e=21.5 \[\frac{|a-10|}{10}\times 100=21.5\] \[|a-10|=2.15\] \[a-10=\pm 2.15\] either a=10-2.15 or a=10+2.15 Do you get this?

OpenStudy (anonymous):

ya that makes sense!! thanks so much!